# Bayesian Nonparametrics for Principal Stratification with Continuous Post-Treatment Variables

### v3

@@ 6 Discussion
In this paper, we introduced the we proposed the Confounders-Aware SHared Atoms Bayesian Mixture Model (CASBAH) to address critical challenges within the principal stratification framework. Our approach leverages the dependent Dirichlet process to define a highly flexible and adaptable model for the potential post-treatment variables. An innovate feature of CASBAH is that we allow the distribution of potential post-treatment variables to share information through the shared-atoms of the Bayesian nonparametric prior. This enables a more accurate identification of the principal strata while relaxing the standard assumptions of the causal framework of the principal stratification.
CASBAH data-adaptively identifies the principal strata and imputes missing post-treatment variables and outcomes without the need to set predefined (and subjective) thresholds on the continuous post-treatment values to identify the principal strata. Moreover, the proposed model allows for the quantification of the uncertainty in membership in the principal strata. Theorem 1 provides identification for the proposed causal estimands, while Theorem 2 illustrates how the probability of a unit belonging to a dissociative stratum is nonzero and depends on its observed covariates.
@@ Appendix A Conjugate prior distribution for probit regression parameters
Appendix B Proof Theorem 1
Theorem 1 in Section 3.2 wants to rewrite the causal estimands in Definition 1 that are not identifiable.
For a clearer notation, we use the notation $\mathbb{E}_{A}\mathbb{E}(B\mid A)$ to indicate $\int_{\mathcal{A}}\mathbb{E}(B\mid A=a)\Pr(A=a)da$ where $\mathcal{A}$ is the support of variable A. The general causal estimand can be rewritten as following
$\mathbb{E}\big\{Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0}\big\}$
$\quad=\mathbb{E}\big\{Y_{i}(1)\mid S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0}\big\}-\mathbb{E}\big\{Y_{i}(0)\mid S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0}\big\}$
$\quad=\mathbb{E}_{X_{i}\mid S_{i}^{(1)},S_{i}^{(0)}}\mathbb{E}\big\{Y_{i}(1)\mid S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big\}$-
$\quad\quad\quad\mathbb{E}_{X_{i}\mid S_{i}^{(1)},S_{i}^{(0)}}\mathbb{E}\big\{Y_{i}(0)\mid S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big\}$
$\quad=\mathbb{E}_{X_{i}\mid S_{i}^{(1)},S_{i}^{(0)}}\mathbb{E}\big\{Y_{i}\mid T_{i}=1,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big\}$-
$\quad\quad\quad\mathbb{E}_{X_{i}\mid S_{i}^{(1)},S_{i}^{(0)}}\mathbb{E}\big\{Y_{i}\mid T_{i}=0,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big\},$
where the first and second equality invokes the properties of expectation and the third equality is due to the no unmeasured confounding and SUTVA assumption. The inner expectation of each integral, for the treatment value $t\in\{0,1\}$, can be decomposed as
$\mathbb{E}\big\{Y_{i}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big\}$
$\quad=\mathbb{E}_{P_{i}(1),P_{i}(0)\mid T_{i},S_{i}^{(1)},S_{i}^{(0)},X_{i}}\mathbb{E}\big\{Y_{i}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big\}$
$\quad=\mathbb{E}_{P_{i}(1),P_{i}(0)\mid T_{i},S_{i}^{(1)},S_{i}^{(0)},X_{i}}\mathbb{E}\big\{Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big\},$
where the second equality invokes the property that the outcome $Y_{i}$ is independent to the latent cluster-allocation $(S_{i}^{(0)},S_{i}^{(1)})$ given the potential post-treatment variable $(P_{i}(0),P_{i}(1))$, by definition of the post-treatment variable model.
Therefore, we can rewritten it explicitly, as reported in Theorem 1,
$\mathbb{E}\big\{Y_{i}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big\}$
$\quad=\int_{p_{0}p_{1}}\mathbb{E}\big\{Y_{i}\mid T_{i}=1,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big\}$
$\quad\quad\times Pr\big(P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big)dp_{0}p_{1}dx$ (14)
The inner expectations $\mathbb{E}\big\{Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big\}$ is estimate with the outcome model $Y_{i}\mid T_{i},X_{i},P_{i}(1),P_{i}(0)$, as well the probability $\Pr\big(P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big)$, that is defined by the model for the potential post-treatment variables.
Appendix C Proof Dissociative Stratum Probability
@@
Appendix D Posterior Computation
@@
Appendix E More simulations details
@@
Appendix F More application details and results

### v4

@@ 6 Discussion
In this paper, we proposed the Confounders-Aware SHared Atoms Bayesian Mixture Model (CASBAH) to address critical challenges within the principal stratification framework. Our approach leverages the dependent Dirichlet process to define a highly flexible and adaptable model for the potential post-treatment variables. An innovative feature of CASBAH is that we allow the distribution of potential post-treatment variables to share information through the treatment levels via the shared-atoms of the Bayesian nonparametric prior. This enables a more accurate identification of the principal strata while relaxing the standard assumptions of the causal framework of the principal stratification.
CASBAH data-adaptively identifies the principal strata and imputes missing post-treatment variables and outcomes without the need to set predefined (and subjective) thresholds on the continuous post-treatment values to identify the principal strata. Moreover, the proposed model allows for the quantification of the uncertainty in membership in the principal strata. Theorem 1 provides identification for the proposed causal estimands, while Theorem 2 illustrates how the probability of a unit belonging to a dissociative stratum is nonzero a priori and depends on its observed covariates.
@@ Appendix A Conjugate prior distribution for probit regression parameters
Appendix B Proof Dissociative Stratum Probability
@@
Appendix C Posterior Computation
@@
Appendix D More simulations details
@@
Appendix E More application details and results